# Guided Transfer Learning for Discrete Diffusion Models

### v1

@@ Abstract
Discrete diffusion models achieve strong performance across language and other discrete domains, providing a powerful alternative to autoregressive models. However, their strong performance relies on large training datasets, which are costly or risky to obtain, especially when adapting to new domains. Transfer learning is the natural way to adapt pretrained discrete diffusion models, but current methods require fine-tuning large diffusion models, which is computationally expensive and often impractical. Building on ratio-based transfer learning for continuous diffusion, we provide Guided Transfer Learning for discrete diffusion models (GTL). This enables sampling from a target distribution without modifying the pretrained denoiser. The same guidance formulation applies to both discrete-time diffusion and continuous-time score-based discrete diffusion, yielding a unified treatment. Guided discrete diffusion often requires many forward passes of the guidance network, which becomes impractical for large vocabularies and long sequences. To address this, we further present an efficient guided sampler that concentrates evaluations on planner-selected positions and top candidate tokens, thus lowering sampling time and computation. This makes guided language modeling practical at scale for large vocabularies and long sequences. We evaluate GTL on sequential data, including synthetic Markov chains and language modeling, and provide empirical analyses of its behavior. This is a pre-print.
@@ 1 Introduction
Diffusion models (ho2020denoisingdiffusionprobabilisticmodels; song2021scorebasedgenerativemodelingstochastic) have become the dominant approach for generative modeling in continuous domains like image, video, and audio (lovelace2023latentdiffusionlanguagegeneration; ho2022videodiffusionmodels; liu2023audioldmtexttoaudiogenerationlatent). More recently, discrete diffusion models for language have achieved competitive performance and, by denoising tokens in parallel and in any order, offer practical advantages for controllable infilling and editing compared to left-to-right autoregressive decoding (gong2025scalingdiffusionlanguagemodels; li2022diffusionlm). Among existing discrete diffusion approaches, masked diffusion models (MDMs) have demonstrated remarkable performance in tasks such as reasoning (nie2025scalingmaskeddiffusionmodels; zheng2025maskeddiffusionmodelssecretly), planning (ye2025autoregressiondiscretediffusioncomplex), and infilling (gong2025scalingdiffusionlanguagemodels), often performing on par with state-of-the-art auto-regressive language models (ALMs).
An open problem in the field is to develop transfer learning approaches that enable leveraging the incredible performance of large pre-trained diffusion models in domains with scarce data (ouyang2024transferlearningdiffusionmodels). In the continuous domain, recent works have explored transfer learning approaches based domain classifiers (ratio estimators) (zhong2025domainguidancesimpletransfer; ouyang2024transferlearningdiffusionmodels). Such methods bypass the need of updating millions of parameters, providing an alternative to adapter-based finetuning (moon2022finetuning; xie2023difffitunlockingtransferabilitylarge) and reinforcement learning (fan2023dpokreinforcementlearningfinetuning; clark2024directlyfinetuningdiffusionmodels) in scenarios of extreme scarcity.
In this paper, we tackle the problem of extending domain classifier-based sampling for discrete diffusion models. We first develop the theoretical analogues of (ouyang2024transferlearningdiffusionmodels), leading to a first naïve implementation of (8) (In supplementary materials, algorithm 2). However, we show that such naïve implementation requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per denoising step, scaling poorly with sequence length $L$ and vocabulary size $|\mathcal{V}|$, limiting its applicability to domains such as language modeling.
Motivated by these challenges, we introduce Guided Transfer Learning (GTL) for discrete diffusion models. Our method leverages a pretrained source discrete diffusion model and introduces a compact ratio network that estimates the density ratio between the target and source data distributions. We demonstrate that sampling from the target domain can be achieved solely by training this ratio-based guidance module, enabling effective target sampling without updating the denoiser. This substantially reduces the training cost compared to conventional fine-tuning approaches. We establish a closed-form solution for the sampling method, and we further extend this argument to continuous-time score-based models. This generalization allows our approach to cover the two dominant diffusion frameworks: score-based models (lou2024discretediffusionmodelingestimating; sun2023scorebasedcontinuoustimediscretediffusion) and discrete-time diffusion models (austin2023structureddenoisingdiffusionmodels), making the method broadly applicable to diffusion modeling. This GTL construction does not follow directly from continuous diffusion models (ouyang2024transferlearningdiffusionmodels), since score-matching diffusion is not defined for categorical discrete tokens and cannot be derived from the score-matching loss function (song2021scorebasedgenerativemodelingstochastic). We further introduce a guided sampling algorithm that reduces the number of guidance calls per denoising step by focusing computation on planner-selected tokens. This design improves efficiency. Previewing our results, GTL surpasses vanilla and fine-tuned diffusion across all data-scarcity regimes while training only  7% as many parameters (Fig. 1).
@@
Figure 1: GTL outperforms with  7% of the parameters. MAUVE ($\uparrow$) vs. fraction of target-domain training data (e.g., 100% = all 79,631 arXiv Physics abstracts; $1\%\approx 800$). Our Guided Transfer Learning (GTL, green) achieves the highest MAUVE across all data regimes while training only a 4.1M parameter ratio model outperforming both Vanilla (trained only on target data) and Fine-tuned (source-pretrained then fine-tuned on target) models. Points show mean, and error bars show standard error over 3×3 (train×sample) seeds. See Section 5.2 for the experimental setup and evaluation details.
2 Related Work
Discrete Diffusion Models
Discrete diffusion models (DDMs) are widely used for biological sequence generation (e.g., DNA, RNA, proteins) as well as language modeling (Sarkar2024; yu2025discretediffusionlargelanguage). Continuous‑noise variants that corrupt embeddings or latents are likely to suffer from embedding‑to‑token projection mismatch  (li2022diffusionlmimprovescontrollabletext; lovelace2023latentdiffusionlanguagegeneration). Beyond that, two widely adopted formulations operate entirely in discrete space, avoiding continuous mappings. First, simplified masked diffusion models such as MLDM and MD4 follow the D3PM framework, using discrete‑time transition matrices with a denoiser trained to invert the forward kernels (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion; austin2023structureddenoisingdiffusionmodels). Second, continuous‑time discrete diffusion estimates a discrete score, which represents the log‑ratio between adjacent‑timestep marginals. This induces a continuous-time Markov chain (CTMC) (lou2024discretediffusionmodelingestimating; meng2023concretescorematchinggeneralized; campbell2022continuoustimeframeworkdiscrete).
Guided Diffusion Models
In both continuous and discrete settings, guidance strategies like classifier‑based or classifier‑free are widely used. In the discrete case, classifier‑free guidance typically performs two forward passes at inference time, one conditional and one unconditional, while training a single large model on a mixture of conditional prompt–input pairs and unconditional data (huang2025ctrldiffboostinglargediffusion). Closer to our approach is classifier‑based guidance, which steers generation using a classifier $p(y\mid x_{t},t)$ while the denoiser is trained on the full corpus; sampling is biased toward the target label $y$ in the data distribution (schiff2025simpleguidancemechanismsdiscrete; nisonoff2025unlockingguidancediscretestatespace). In both discrete- and continuous-time formulations, this typically yields per-token factorized guidance, which incurs many computational steps, scaling with $\mathcal{O}(L|\mathcal{V}|)$ evaluations per denoising step. To mitigate this cost, (nisonoff2025unlockingguidancediscretestatespace) proposes a first-order Taylor approximation that reduces guidance evaluations to a forward and backward pass. TESS2 (tae2025tess2largescalegeneralist) instead applies gradient ascent on a reward with respect to the model logits during inference to improve response quality. Finally, the Energy‑based Diffusion Language Model (xu2025energybaseddiffusionlanguagemodels) defines an energy over predicted clean sequences and uses it to importance‑weight the denoising step, reducing the train–sample mismatch in discrete diffusion models.
Transfer Learning
Fine-tuning and transfer learning share the goal of adapting a model to a new task or domain. Fine-tuning is computationally expensive because it updates the pretrained denoiser’s weights on new data. Prior work on continuous diffusion models explores reinforcement learning (fan2023dpokreinforcementlearningfinetuning; clark2024directlyfinetuningdiffusionmodels) and adapter-based fine-tuning (moon2022finetuning; xie2023difffitunlockingtransferabilitylarge). In contrast, transfer learning for diffusion remains comparatively underexplored. Transfer Learning for Diffusion Models (TLDM) proposes a ratio-estimation technique for continuous score-based models, but demonstrates effectiveness on synthetic data and a one-dimensional heartbeat dataset, leaving applications to higher-dimensional and large-scale domains open.
@@ 3 Background
Discrete Denoising Diffusion Probabilistic Models (D3PMs) (austin2023structureddenoisingdiffusionmodels) provide a flexible framework for diffusion over discrete state spaces. The forward Markov chain is $p(\mathbf{z}_{t}\mid\mathbf{z}_{s})=\text{Cat}\!\left(\mathbf{z}_{t};\,Q_{t\mid s}\mathbf{z}_{s}\right)$ where $Q_{t\mid s}\in\mathbb{R}^{N\times N}$ is a transition matrix whose $(i,j)$ entry is the probability of moving to state $i$ at time $t$ from state $j$ at time $s$. To fit the generative model $p_{\theta}$ to the data distribution over $\mathbf{x}_{0}$, we minimize the negative ELBO (NELBO). Different choices of $Q_{t\mid s}$ induce different corruption behaviors, e.g., uniform, absorbing, or discretized-Gaussian transitions.
@@
The absorbing-state variant or often called masked diffusion, replaces tokens with a special [MASK] token (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion). It can be viewed as an interpolation between the clean distribution and a noise prior over the vocabulary. For a clean token $\mathbf{x}_{0}$, the forward transition at time $t$ is
@@
$p(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}_{0})$
$=\begin{cases}\operatorname{Cat}\!\left(\mathbf{z}_{s};\,\mathbf{z}_{t}\right),&\text{if }\mathbf{z}_{t}\neq\mathbf{m},\\[6.0pt] \operatorname{Cat}\!\left(\mathbf{z}_{s};\,\dfrac{(1-\alpha_{s})\mathbf{m}+(\alpha_{s}-\alpha_{t})\mathbf{x}_{0}}{1-\alpha_{t}}\right),&\text{if }\mathbf{z}_{t}=\mathbf{m}.\end{cases}$ (2)
Hence, when a token is unmasked ($\mathbf{x}_{t}\neq\mathbf{m}$) the step is deterministic: $\mathbf{x}_{s}=\mathbf{x}_{t}$. Otherwise, the posterior is a linear interpolation between the mask distribution and the clean token. This structure yields the simplified diffusion loss (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion):
$\mathcal{L}_{\text{diffusion}}$ $=\sum_{i=1}^{T}D_{\mathrm{KL}}\!\left(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\right)$ (3)
$=\sum_{i=1}^{T}\mathbb{E}_{q}\!\left[\frac{\alpha_{t(i)}-\alpha_{s(i)}}{1-\alpha_{t(i)}}\log\big\langle\mathbf{x}_{\theta}(\mathbf{z}_{t(i)}),\mathbf{x}_{0}\big\rangle\right],$ (4)
@@
Diffusion models provide strong controllability through both classifier-based and classifier-free guidance. These two approaches correspond to different ways of expressing the backward generative process conditioned on $y$. In this work, we focus on the classifier-based variant, which allows us to rearrange the conditional generative process using Bayes’ rule as $p^{\gamma}(\mathbf{z}_{s}\mid\mathbf{z}_{t},y)=p(y\mid\mathbf{z}_{s},\mathbf{z}_{t})^{\gamma}\,p(\mathbf{z}_{s}\mid\mathbf{z}_{t})/Z$, with $Z$ being a normalizing constant.
@@
$p^{\gamma}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)},y\right)$
$=\prod_{\ell=1}^{L}\frac{p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}p\!\left(y\mid\tilde{\mathbf{z}}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}$ (5)
Here, $\mathcal{V}$ denotes the vocabulary set, and $p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)$ represents the classifier probability given the sequence $\mathbf{z}_{t}^{(1:L)}$ with the token $\mathbf{z}_{s}$ placed at position $\ell$. However, this formulation comes with the drawback of requiring $\mathcal{O}(L\cdot|\mathcal{V}|)$ forward passes through the classifier model.
@@ 4 Guided Transfer Learning for DDMs
We propose Guided Transfer Learning for Discrete Diffusion Models, a framework that enables transfer learning for discrete diffusion models. We first formalize the problem setup, then present the mathematical formulation of our method, and finally introduce a new sampling algorithm that leverages our results to make transfer learning broadly applicable across datasets.
Given the problem statement and access to both domains, one could fine-tune the source network $p_{\theta}$ by maximizing (4), thereby optimizing the entire parameter set. For large language models, this is often costly and slow due to the scale of parameters. Moreover, when the target domain provides far fewer samples, fine-tuning alone rarely yields a strong generative model in terms of diversity and quality.
Our method is based on the following theorem, which proposes a different approach to this problem: it leverages a pretrained source-domain diffusion model together with a ratio estimator between the two domains. We assume the two diffusion models share the same forward process $p(\mathbf{z}_{t}\mid\mathbf{x}_{0})=q(\mathbf{z}_{t}\mid\mathbf{x}_{0})$ for all $t$, that the density ratio $q(\mathbf{x}_{0})/p(\mathbf{x}_{0})$ is well defined.
Relation to TLDM. TLDM (ouyang2024transferlearningdiffusionmodels) proves a transfer result for continuous, score-based diffusion and is not directly applicable to discrete state spaces. The theorem is stated in $\mathbb{R}^{d}$ and relies on score matching with differentiable densities, requiring gradients such as $\nabla_{x_{t}}\log q_{t}(x_{t})$. In the discrete setting (categorical tokens), such Euclidean scores are undefined. Beyond this, the modeling and training objectives differ: continuous diffusion injects Gaussian noise and optimizes score-matching Mean Squared Error, whereas discrete diffusion uses categorical transition matrices (e.g., absorbing/uniform) and is trained by minimizing KL between categorical true posteriors $q(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}_{0})$ and $q_{\psi}(\mathbf{z}_{s}\mid\mathbf{z}_{t})$.
@@
In (7), $p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\approx p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ is the trained denoiser on the source domain, and the expectation factor $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\tfrac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right]$ is the guidance ratio. Thus, the target diffusion process can be approximated with two networks $\psi^{\star}=[\theta,\phi]$. The proof of this theorem can be found in the supplementary materials. The advantage of this approach is that it does not require fine-tuning the pretrained discrete diffusion model. Instead, the main computational cost is shifted to training a domain classifier using both domains as data. Importantly, the domain classifier can be much smaller than the discrete diffusion model. Moreover, density ratio estimation provides more flexibility in regularization when dealing with highly imbalanced data. It only requires distinguishing domains rather than modeling the full target distribution, which makes the training dynamics between the discriminative and generative models more stable.
The result in (ouyang2024transferlearningdiffusionmodels) is derived only for score-based diffusion in continuous data domains. But our formulation in (7) offers two advantages: (i) it permits flexible definitions of the forward and reverse discrete diffusion processes, and (ii) the guidance term can be trained under arbitrary noise processes, in either continuous- or discrete-time settings.
In the supplementary materials (Theorem 3), we extend the result to the score-based, continuous-time diffusion setting over discrete state spaces. The key mathematical step is to trace back through to the ELBO formulation used in (6). For the masked (absorbing-state) diffusion training objective, this link is directly shown in Equation (4).
Extension to Sequence
Leveraging our theoretical results, (7) closely parallels (5). Instead of a classifier $p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)$, we use a ratio network $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\tfrac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right]$. To extend this to sequences, define $\mathbf{z}_{t,s^{\ell}}^{(1:L)}\coloneqq[\,\mathbf{z}_{t}^{(1:\ell-1)},\,\mathbf{z}_{s}^{(\ell)},\,\mathbf{z}_{t}^{(\ell+1:L)}\,]$, i.e., the sequence identical to $\mathbf{z}_{t}^{(1:L)}$ except at position $\ell$, where the token is replaced by $\mathbf{z}_{s}^{(\ell)}$. This yields the following factorized, ratio-guided conditional (with guidance strength $\gamma\geq 0$):
$q_{\theta,\phi}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)}\right)$

### v2

@@ Abstract
Discrete diffusion models (DMs) have achieved strong performance in language and other discrete domains, offering a compelling alternative to autoregressive modeling. Yet this performance typically depends on large training datasets, challenging the performance of DMs in small-data regimes—common under real-world constraints. Aimed at this challenge, recent work in continuous DMs suggests that transfer learning via classifier ratio–based guidance can adapt a pretrained DM to a related target distribution, often outperforming alternatives such as full-weight fine-tuning on the target data. By contrast, transfer learning for discrete DMs remains unexplored. We address this gap by exploring practical analogues of ratio-based transfer learning for discrete DMs. Our theoretical analysis shows that a direct extension of existing ratio-based guidance is computationally prohibitive, scaling with vocabulary size. To overcome this limitation, we introduce a scheduling mechanism that yields a practical algorithm, Guided Transfer Learning for discrete diffusion models (GTL). GTL enables sampling from a target distribution without modifying the pretrained denoiser and reduces the cost to linear scaling in vocabulary size, which in turn supports longer sequence generation. We evaluate GTL on sequential data, including synthetic Markov chains and language modeling tasks, and provide a detailed empirical analysis of its behavior. The results highlight a clear trade-off: when target datasets are large, weight fine-tuning is often preferable, whereas GTL becomes increasingly effective as target data shrinks. Finally, we experimentally demonstrate a key failure mode of GTL: when the source and target distributions overlap poorly, the ratio-based classifier required for guidance becomes unreliable, limiting transfer performance.
@@ 1 Introduction
Diffusion models (DMs) (Ho et al., 2020; Song et al., 2021) have become the dominant approach for generative modeling in continuous domains like image, video, and audio (Lovelace et al., 2023; Ho et al., 2022; Liu et al., 2023). More recently, discrete DMs for language have achieved competitive performance and, by denoising tokens in parallel and in any order, offer practical advantages for controllable infilling and editing compared to left-to-right autoregressive decoding (Gong et al., 2025; Li et al., 2022a). Among existing discrete diffusion approaches, masked diffusion models (MDMs) have demonstrated remarkable performance in tasks such as reasoning (Nie et al., 2025; Zheng et al., 2025), planning (Ye et al., 2025), and infilling (Gong et al., 2025), often performing on par with state-of-the-art auto-regressive language models (ALMs).
Training DMs require large amounts of data (Wang et al., 2023). And it remains an open question how to best leverage their powerful generative modeling capabilities in domains where scarse data is available (Zhang et al., 2025), a common scenario in applications like medical imaging, where privacy and limited public availability constrain dataset size (Xie et al., 2025). This problem motivates the exploration of transfer learning techniques, where the goal is leverage a DM pretrained on a large dataset for generative modeling in a related domain where data is more scarse. One possible solution is to fine-tune the pretrained DM’s denoiser on the target data (Xie et al., 2023b), but this becomes ineffective in very small data regimes (Moon et al., 2022). To address this, recent works have proposed transfer learning to tackle the small data regime by using classifier-based guidance (CBG) (Zhong et al., 2025; Ouyang et al., 2024), which estimates a density-ratio between the source and target domain and uses it as a diffusion guidance mechanism. Such methods are more effective than finetuning in scenarios of extreme scarcity, as they bypass the need of updating millions of parameters (Moon et al., 2022; Xie et al., 2023a). However, these methods have only been investigated in continuous domains (e.g., images). By contrast, transfer learning for discrete DMs such as language models, remains largely unexplored.
Outside of transfer learning, Schiff et al. (2025); Nisonoff et al. (2025) have explored CBG mechanisms for discrete DMs. These methods apply mainly to domains with small $|\mathcal{V}|$ and modest $L$, such as biological sequences, leaving scalable CBG for language modeling as an open problem. This constaints are due to the necessity of token substitutions across all positions and vocabulary items during score computation. More precisely, letting $L$ be the sequence length and $\mathcal{V}$ the vocabulary, naive CBG requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per step, which is infeasible for long sequences and large vocabularies. In this paper, we aim to leverage CBGs for transfer learning, while overcoming the limiting small vocabulary sizes.
First, we present a theoretical analysis deriving a closed-form ratio-guided reverse transition for discrete-time diffusion and extend this principle to continuous-time, score-based discrete diffusion, covering both score-based models (Lou et al., 2024; Sun et al., 2023) and discrete-time diffusion models (Austin et al., 2023). Such derivation is architecture- and corruption-agnostic—deriving the result from the evidence lower bound (ELBO). Next, we address the scaling problem of CBGs for discrete DMs. For this purpose, we introduce an efficient sampler that makes guided language modeling for masked discrete DMs practical by avoiding $\mathcal{O}(L|\mathcal{V}|)$ guidance calls, concentrating ratio evaluations on planner-selected positions and a top-$n_{\text{ratio}}$ candidate set.
[…]
Leveraging this result, we introduce a method termed Guided Transfer Learning (GTL). The core idea of GTL is to freeze a pretrained source denoiser and learn a lightweight ratio network from mixed source/target data to estimate target-to-source density ratios; at sampling time, we use this ratio with a learned planner network as guidance to reweight the denoiser’s reverse transition and generate target-domain samples without fine-tuning. This procedure allows to effectively utilize transfer learning based on CBG for discrete DMs. Previewing our results, GTL outperforms vanilla and fine-tuned diffusion across data-scarcity regimes while training only $\sim 7\%$ as many parameters (Fig. 1). In a setting common in practice—where the pretraining corpus already contains some in-domain target samples—we find that GTL performs better and yields more stable performance. Our result also reveals limitations of CBG for discrete modeling: under weak domain overlap, the interaction between guidance strength $\gamma$ and candidate pruning can amplify ratio errors. In this regime, larger $\gamma$ or small $n_{\text{ratio}}$ may destabilize denoising and hurt quality.
@@ 2 Background
Discrete Denoising Diffusion Probabilistic Models (D3PMs) (Austin et al., 2023) provide a flexible framework for diffusion over discrete state spaces. The forward Markov chain is $p(\mathbf{z}_{t}\mid\mathbf{z}_{s})=\text{Cat}\!\left(\mathbf{z}_{t};\,Q_{t\mid s}\mathbf{z}_{s}\right)$ where $Q_{t\mid s}\in\mathbb{R}^{N\times N}$ is a transition matrix whose $(i,j)$ entry is the probability of moving to state $i$ at time $t$ from state $j$ at time $s$. To fit the generative model $p_{\theta}$ to the data distribution over $\mathbf{x}_{0}$, we minimize the negative ELBO (NELBO). Different choices of $Q_{t\mid s}$ induce different corruption behaviors, e.g., uniform, absorbing, or discretized-Gaussian transitions.
@@
The absorbing-state variant or often called masked diffusion, replaces tokens with a special [MASK] token (Sahoo et al., 2024; Shi et al., 2025). It can be viewed as an interpolation between the clean distribution and a noise prior over the vocabulary. For a clean token $\mathbf{x}_{0}$, the forward transition at time $t$ is
@@
$p(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}_{0})=\begin{cases}\operatorname{Cat}\!\left(\mathbf{z}_{s};\,\mathbf{z}_{t}\right),&\text{if }\mathbf{z}_{t}\neq\mathbf{m},\\[6.0pt] \operatorname{Cat}\!\left(\mathbf{z}_{s};\,\dfrac{(1-\alpha_{s})\mathbf{m}+(\alpha_{s}-\alpha_{t})\mathbf{x}_{0}}{1-\alpha_{t}}\right),&\text{if }\mathbf{z}_{t}=\mathbf{m}.\end{cases}$ (2)
Hence, when a token is unmasked ($\mathbf{x}_{t}\neq\mathbf{m}$) the step is deterministic: $\mathbf{x}_{s}=\mathbf{x}_{t}$. Otherwise, the posterior is a linear interpolation between the mask distribution and the clean token. This structure yields the simplified diffusion loss (Sahoo et al., 2024; Shi et al., 2025):
$\mathcal{L}_{\text{diffusion}}$ $=\sum_{i=1}^{T}D_{\mathrm{KL}}\!\left(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\right)=\sum_{i=1}^{T}\mathbb{E}_{q}\!\left[\frac{\alpha_{t(i)}-\alpha_{s(i)}}{1-\alpha_{t(i)}}\log\big\langle\mathbf{x}_{\theta}(\mathbf{z}_{t(i)}),\mathbf{x}_{0}\big\rangle\right],$ (3)
@@
Diffusion models provide strong controllability through both classifier-based and classifier-free guidance. These two approaches correspond to different ways of expressing the backward generative process conditioned on $y$. In this work, we focus on the classifier-based variant, which allows us to rearrange the conditional generative process using Bayes’ rule as $p^{\gamma}(\mathbf{z}_{s}\mid\mathbf{z}_{t},y)=p(y\mid\mathbf{z}_{s},\mathbf{z}_{t})^{\gamma}\,p(\mathbf{z}_{s}\mid\mathbf{z}_{t})/Z$, with $Z$ being a normalizing constant. (Schiff et al., 2025; Nisonoff et al., 2025)
@@
$p^{\gamma}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)},y\right)=\prod_{\ell=1}^{L}\frac{p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}p\!\left(y\mid\tilde{\mathbf{z}}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}$ (4)
Here, $\mathcal{V}$ denotes the vocabulary set, and $p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)$ represents the classifier probability given the sequence $\mathbf{z}_{t}^{(1:L)}$ with the token $\mathbf{z}_{s}$ placed at position $\ell$. However, this formulation comes with the drawback of requiring $\mathcal{O}(L\cdot|\mathcal{V}|)$ forward passes through the classifier model, which is infeasible for long sequences and large vocabularies. Because each evaluation corresponds to a distinct single-token intervention on the full sequence, this cost is not removed by straightforward parallelization.
@@ 3 Guided Transfer Learning for DDMs
We introduce Guided Transfer Learning (GTL), a transfer framework for discrete diffusion models that adapts a pretrained source model to a target domain without updating the denoiser. Our approach is orthogonal to fine-tuning: rather than updating the denoiser, GTL adapts sampling by reweighting the source reverse transitions with a learned density-ratio guidance model. In this section, we first present the core transfer identity for a single denoising step, then extend it to sequences and develop an efficient guided sampler.
Why not fine-tune? Fine-tuning adapts the source denoiser $p_{\theta}$ to the target domain by optimizing the diffusion objective (e.g., equation 3) on target samples. This optimization becomes expensive for large language models and tends to overfit when target data are scarce, often degrading diversity. These small-data failures motivate our approach: rather than updating millions of denoiser parameters, we reuse $p_{\theta}$ as-is and place the adaptation burden on a much smaller auxiliary model.
Key Assumption. This auxiliary model is meaningful only if the two domains share the same corruption mechanism. Let $p_{X}$ denote the source data distribution and $q_{X}$ the target data distribution over $\mathbf{x}_{0}$. We assume a shared forward process, $p(\mathbf{z}_{t}\mid\mathbf{x}_{0})=q(\mathbf{z}_{t}\mid\mathbf{x}_{0})$ for all $t$, which holds whenever both domains use the same discrete forward kernel (e.g., an identical masking schedule or transition family). Under this shared corruption, differences between the domains can be expressed at the data level through the density ratio $r(\mathbf{x}_{0})\coloneqq q_{X}(\mathbf{x}_{0})/p_{X}(\mathbf{x}_{0}),$ assumed well-defined on the support of interest. This ratio quantifies how much more (or less) likely a clean sequence is under the target distribution than under the source, and it is exactly the quantity needed to reweight source reverse transitions into target reverse transitions.
The next theorem makes this reweighting precise for a single reverse transition: it shows that the reverse kernel minimizing the target KL objective can be written as the source reverse kernel multiplied by a ratio-dependent weight.
@@
Practical implementation. Equation 6 (proved in Appendix C) expresses the target reverse transition as a reweighted version of the source reverse transition. This expression suggests a direct implementation with two components: a frozen source denoiser that provides $p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ and a learned ratio module that provides the ratio-dependent weight. Concretely, we approximate the source kernel with the pretrained denoiser $p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$, and we learn a ratio function $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right].$ Together, these two networks define the GTL sampler, $\psi^{\star}=[\theta,\phi]$, which generates target-domain samples while keeping the denoiser fixed.
Why ratio guidance helps in the low-data regime? The key advantage of equation 6 is that it eliminates the need to fine-tune the denoiser on scarce target data. Instead, adaptation is shifted to learning a domain discriminator (or an equivalent ratio estimator) from mixed source/target samples, which can be orders of magnitude smaller than the diffusion model. This shift is especially helpful under strong source–target imbalance: ratio estimation only needs to separate domains, rather than model the full target distribution, and it admits flexible regularization that tends to stabilize training in data-scarce settings.
Relation to TLDM and generality of our formulation. TLDM (Ouyang et al., 2024) derives a ratio-transfer rule only for continuous data score-based diffusion in $\mathbb{R}^{d}$, where guidance is expressed through Euclidean score gradients such as $\nabla_{x_{t}}\log q_{t}(x_{t})$. These gradients are not available for categorical tokens, and discrete diffusion is commonly trained by minimizing KL divergences between categorical posteriors rather than score-matching MSE. Our derivation therefore works directly with discrete forward and reverse kernels and the ELBO-style objective in equation 5. This formulation is architecture- and corruption-agnostic: it applies to a range of discrete forward processes (e.g., uniform, absorbing/masked, and general D3PM transitions) and does not depend on a particular parameterization of the denoiser in benefit to TLDM.
Continuous-Time GTL for Discrete Diffusion. Because the argument is expressed at the level of the ELBO/KL objective, the same ratio-transfer principle carries over to continuous-time discrete diffusion whenever the training loss can be related back to an ELBO (e.g., equation 3 and equation 36). In Appendix (Theorem 3), we instantiate this transfer rule for continuous-time, score-based diffusion over discrete state spaces by reweighting the reverse-rate matrix with the learned ratio.
Sequence Formulation.
We now apply Theorem 1 to length-$L$ token sequences. As in prior work on discrete classifier guidance, we adopt a token-factorized reverse transition, which yields a practical sampler while keeping the denoiser unchanged. Concretely, we replace the classifier term $p(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)})$ in equation 4 with a learned ratio model. In the sequence setting, the ratio model is evaluated on a single-token intervention of the current noisy sequence. Define $\mathbf{z}_{t,s^{\ell}}^{(1:L)}\coloneqq[\,\mathbf{z}_{t}^{(1:\ell-1)},\,\mathbf{z}_{s}^{(\ell)},\,\mathbf{z}_{t}^{(\ell+1:L)}\,]$, i.e., the sequence obtained by replacing position $\ell$ in $\mathbf{z}_{t}^{(1:L)}$ by a candidate token $\mathbf{z}_{s}^{(\ell)}$. With guidance strength $\gamma\geq 0$, the resulting ratio-guided, factorized reverse transition is equation 7.